OpenStudy (hpfan101):

Find y' and y'' of: \[y=\cos(\sin3x)\] I solved for y' and not sure how to go about y'' \[y=\cos(\sin3x) \rightarrow y'=-\cos(3x)\times \sin(\sin3x) \]

OpenStudy (hpfan101):

y'=-3cos(3x) (sin(sin3x) Forgot the 3

zepdrix (zepdrix):

\[\large\rm y'=(-3\cos3x)\sin(\sin3x)\]Ok your first derivative looks great. Let's first "set up" our second derivative,\[\large\rm y''=\color{royalblue}{(-3\cos3x)'}\sin(\sin3x)+(-3\cos3x)\color{royalblue}{(\sin(\sin3x))'}\]

OpenStudy (hpfan101):

Oh I forgot about the product rule. So for the first part, would it be 9sin3x times sin(sin3x)

zepdrix (zepdrix):

\[\large\rm y''=\color{orangered}{(9\sin3x)}\sin(\sin3x)+(-3\cos3x)\color{royalblue}{(\sin(\sin3x))'}\]Sounds right so far!

OpenStudy (hpfan101):

\[9\sin3x \times \sin(\sin3x) + -3\cos3x \times 3\cos(\cos3x)\]

OpenStudy (hpfan101):

I think the second part might be missing something

zepdrix (zepdrix):

\[\large\rm \frac{d}{dx}\sin(\sin3x)=\cos(\sin3x)\cdot\color{royalblue}{(\sin3x)'}\]

zepdrix (zepdrix):

\[\large\rm \frac{d}{dx}\sin(\sin3x)=\cos(\sin3x)\cdot\color{orangered}{(3\cos3x)}\]

OpenStudy (hpfan101):

Ah okay.

OpenStudy (hpfan101):

So our final answer would be: \[y"=(9\sin3x)(\sin(\sin3x))+(-3\cos3x)(\cos(\sin3x)(3\cos3x)\]

zepdrix (zepdrix):

\[\large\rm y''=\color{orangered}{(9\sin3x)}\sin(\sin3x)+(-3\cos3x)\color{royalblue}{(\sin(\sin3x))'}\]\[\large\rm y''=\color{orangered}{(9\sin3x)}\sin(\sin3x)+(-3\cos3x)\color{orangered}{(\cos(\sin3x)\cdot3\cos3x)}\]Mmm good good. Gonna clean it up a little bit? :) heh

zepdrix (zepdrix):

Oh I guess you can't do very much simplification actually..

OpenStudy (hpfan101):

Oh okay. That's fine by me if we can't simplify it :D

OpenStudy (hpfan101):

Thanks for the help!

zepdrix (zepdrix):

Bring the 3's together in the second term, and bring the cos3x's together as well.\[\large\rm y''=9\sin3x\cdot \sin(\sin3x)-9\cos^23x\cdot\cos(\sin3x)\]That looks a little nicer, not a big deal though. np!